# Stress-Testing Long-Context Language Models with Lifelong ICL and Task Haystack

### v1

@@ Appendix C NIAH-style Visualizations
Each figure contains three subfigures: On the left side, we illustrate the results of the original needle-in-a-haystack evaluation [Kamradt, 2023]. We use the question “What is the best thing to do in San Francisco?” and the needle “The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day.” The score is the token-level recall of the model’s response. In the middle, we visualize the results of the Scale-Shot setting. On the right side, we visualize the results of the Scale-Task setting. See §4 for the details of the two scaling settings.
@@
Figure 11: Task Haystack Results on Llama2-7B (32k).
Figure 12: Task Haystack Results on Llama2-7B (80k).
Figure 13: Task Haystack Results on Llama3-8B (1048k).
Figure 14: Task Haystack Results on Llama3-70B (1048k).
@@
Appendix D Fine-grained Diagnostic Reports
@@
D.1 How to interpret the diagnostic report?
@@ D.2 Main Findings
D.3 Visualizations
D.3.1 Mistral-7B, N-task=16, N-shot=8
@@
D.3.2 FILM-7B, N-task=16, N-shot=8
@@
D.3.3 GPT-3.5-Turbo, N-task=16, N-shot=4
@@ D.3.4 GPT-4o, N-task=16, N-shot=8
D.3.5 Mistral-7B, 32-task, 2-shot
@@
D.3.6 Mistral-7B, 64-task, 2-shot

### v2

@@ B.5 Original NIAH Experiments
We experiment with the original needle-in-a-haystack evaluation [Kamradt, 2023] to provide reference. We use the question “What is the best thing to do in San Francisco?” and the needle “The best thing to do in San Francisco is eat a sandwich and sit in Dolores Park on a sunny day.” The metric is the token-level recall of the model’s response. We report the pass rates in Table 12 and visualize the results in the first column of figures in Fig. 9-20.
B.6 Interleaving Examples from Multiple Tasks
The complexity of Lifelong ICL can be further increased by interleaving in-context learning examples from multiple tasks, analogous to a multi-needle NIAH challenge [Hsieh et al., 2024]. We conducted preliminary experiments of this setting, where each ICL example is paired with its task instruction before itself, and ICL examples of different tasks are streamed in a random order. The results are presented in Table 12. We observe that accuracies in this setting (M-acc) are generally lower than those in the non-interleaving Lifelong ICL setting (L-acc), confirming that interleaving examples adds more challenges in locating relevant context. We leave further investigation as future work.
Table 11: Results of the original NIAH evaluation (§B.5).
Model Pass (%) Model Pass (%)
Mistral-7B (32k) 95.3 Llama-3-8B (1048k) 100.0
FILM-7B (32k) 100.0 Yi-6B (200k) 100.0
Llama-2-7B (32k) 95.3 Yi-9B (200k) 100.0
Llama-2-7B (80k) 100.0 Yi-34B (200k) 100.0
Table 12: Results of Interleaving Examples Across Tasks (§B.6). Experiments done in the 16-Task, 4-shot setting. “M-acc” indicates results of interleaving examples.
Model S-acc L-acc M-acc
Mistral-7B (32k) 78.6 74.8 72.1
FILM-7B (32k) 79.6 75.4 74.7
Appendix C Additional Discussion
Table 13: Results of RAG Baseline (§C). Experiments done with 16-Task, 4-Shot. “RAG-acc” indicates the RAG baseline results. Single-task ICL (S-acc) can be seen as an oracle setting with perfect retrieval accuracy.
Model S-acc L-acc RAG-acc
Mistral-7B (32k) 78.6 74.8 79.0
FILM-7B (32k) 79.6 75.4 80.1
Task Haystack can be solved by a RAG baseline.
To examine whether Task Haystack can be addressed by retrieval-augmented generation (RAG) methods, we implemented a simple RAG baseline. We used an off-the-shelf retriever [Zhang, 2024] to select one prompt from all task-specific ICL prompts (i.e., $p_{a_{1}},p_{a_{2}},...,p_{a_{n}}$ as defined in §3). The selected prompt was then prepended before the instruction of the test task. We report the results in Table 13. As expected, Task Haystack, being a retrieval-style task, can be solved by this RAG baseline.
Task Haystack is still meaningful for long-context LM evaluation.
Although Task Haystack can be solved by a RAG baseline, we believe Task Haystack—and retrieval-style tasks more broadly—remain valuable for evaluating long-context models. (1) One advantage of retrieval-style tasks is that it’s much more controllable for ablations. This allows us to carefully investigate whether these long-context LMs behave robustly and as expected. (2) As pointed out by Lee et al. [2024], long-context models have certain benefits over RAG methods, including having a simpler pipeline, better handling of multi-hop queries and mitigating cascading errors. HELMET [Yen et al., 2024], a recently-released long-context benchmark, also incorporates retrieval-style and retrieval-augmented generation tasks.
Task Haystack adds to the axis of contextual understanding in long-context LM evaluation.
Goldman et al. [2024] introduce a taxonomy of long-context LM evaluation with two axes: (1) diffusion: how hard it is to find and extract the necessary information, and (2) scope: how long the necessary information is. In the view of this taxonomy, Task Haystack is having low diffusion (it’s not hard to find relevant information) and small scope (the relevant information is short, only a few ICL examples), similar to the original NIAH. However, Task Haystack is also more challenging than the original NIAH partly due to requiring contextual understanding (or “implicit aggregations” as briefly mentioned in Goldman et al. [2024]) of the context. We believe it may be helpful to add a third axis of “contextual understanding” or “implicit aggregation” to the taxonomy, and Task Haystack can be seen as making progress in this third axis.
Appendix D NIAH-style Visualizations
In Fig. 9-20, we present detailed Task Haystack results for ten open-weight models. In Fig. 20-22 we present results for Gemini-1.5-Flash, GPT-3.5-Turbo and GPT-4o.
Fig. 9-20 each contains three subfigures: On the left side, we illustrate the results of the original needle-in-a-haystack evaluation [Kamradt, 2023], described in §B.5. In the middle, we visualize the results of the Scale-Shot setting. On the right side, we visualize the results of the Scale-Task setting. See §4 for the details of the two scaling settings.
@@ Appendix D NIAH-style Visualizations
Figure 11: Task Haystack Results on Llama-2-7B (32k).
Figure 12: Task Haystack Results on Llama-2-7B (80k).
Figure 13: Task Haystack Results on Llama-3-8B (1048k).
Figure 14: Task Haystack Results on Llama-3-70B (1048k).
@@
Figure 19: Task Haystack Results on Llama-3.1-70B (128k).
Figure 20: Task Haystack Results on Gemini-1.5-Flash (128k).
@@
Appendix E Fine-grained Diagnostic Reports
@@
E.1 How to interpret the diagnostic report?
@@ E.2 Main Findings
E.3 Visualizations
E.3.1 Mistral-7B, N-task=16, N-shot=8
@@
E.3.2 FILM-7B, N-task=16, N-shot=8
@@
E.3.3 GPT-3.5-Turbo, N-task=16, N-shot=4
@@ E.3.4 GPT-4o, N-task=16, N-shot=8
E.3.5 Mistral-7B, 32-task, 2-shot
@@
E.3.6 Mistral-7B, 64-task, 2-shot